OpenStudy (anonymous):

sally wants $7,650 in 2 years using an interest rate of 12% compounded monthly, How much does she need to invest today for this to happen?

OpenStudy (anonymous):

You will need to use this formula where A is final amount (7,650), P is principle amount (what you are solving for), r is annual interest rate (0.12), n is the number of times the interest is compounded yearly (12, since there are 12 months in a year), and t is number of years (2 years): \[ A=P (1 + r/n)^T\] where T=n*t. So solving for P, we have P= A/((1+r/n)^(n*t)) =7650/((1+.12/12)^(12*2)) =6,024.88 So Sally will need to invest $6,024.88 today. Does this make sense? The original formula should help you if you have problems similar to this one.